translate into a algebraic equations :
1) A girl buys three apples and seven oranges for $4.56. If an orange costs $0.28 more than an apple, how much does each fruit cost?

Answer by checkley77(12844)

3A+7O=4.56
O=A+.28
3A+7(A+.28)=4.56
3A+7A+1.96=4.56
10A=4.56-1.96
10A=2.60
A=2.60/10
A=.26 PRICE OF AN APPLE.
O=.26+.28
O=.54 PRICE OF AN ORANGE.
PROOF:
3*.26+7*.54=4.56
.78+3.78=4.56
4.56=4.56